# Ideal Triangulations and Once-Punctured Surface Bundles

### v1

@@ 2.2. Fibers and lw-Taut Surfaces
Let $F$ be a compact, orientable, once-punctured surface with $h\in\text{Aut}(F)$. Let $M_{h}$ be the mapping torus $M_{h}=F\times[0,1]/\sim$, where $(x,0)\sim(h(x),1)$. We say a 3-manifold, $M$, fibers over the circle $S^{1}$ with fiber $F$ if $M$ is homeomorphic to $M_{h}$ for some $h\in\text{Aut}(F)$. This gives a proper embedding of $F$ in $M$ such that the image is incompressible and $\partial$-incompressible. The identification of $M$ with the mapping torus has strong implications on the homology groups $H_{2}(M)$ and $H_{2}(M,\partial M)$. In particular, any incompressible surface properly embedded in $M$ that is homologous to $F$ is homotopic into $F$ [35]. In the case of knots in rational homology 3-spheres this relation is stronger.
@@
An oriented, incompressible, $\partial$-incompressible surface, $S$, properly embedded in $M$ is said to be taut if the homology class $[S]\in H_{2}(M,\partial M)$ is non-trivial, there is no homologically trivial union of components of $S$, and $S$ is $\chi_{-}$-minimizing over all properly embedded surfaces representing $[S]$. If for a triangulation $\mathcal{T}$, $S$ transverse to the 2-skeleton of $\mathcal{T}$, we define the weight of $S$ by $wt(S)=|S\cap\mathcal{T}^{(0)}|$. We say $S$ is lw-taut if it has minimal weight among all taut surfaces representing the homology class $[S]$. Define $kS$ to be the disjoint union of $k$ copies of $S$. If $S$ is lw-taut it follows that $kS$ is an lw-taut surface with homology class $k[S]$. The following from [44] gives existence of normal surfaces that are lw-taut:
@@
For a normal surface $S$, we say the carrier, $C_{S}$ is taut (lw-taut) if every surface carried by $C_{S}$ is taut (lw-taut). The study of faces of $\mathcal{P}_{\mathcal{T}}$ that carry only minimal weight normal surfaces began with Jaco and Oertel [20] which led to an algorithm to detect if a 3-manifold was Haken (contains an essential surface of genus $\geq$ 1). Again from [44] we have the existence of such faces:
@@ 3.1. Crushing Triangulations along Normal Surfaces
We say a 1- or 2-cell of $\mathcal{C}_{X}$ is horizontal if their interior is in $S$ and vertical otherwise. The quadrilateral vertical 2-cells are called trapezoids. There are two trapezoids in a truncated-prism, three in a triangular block, and four in a rectangular block. All other vertical 2-cells are hexagons found in truncated-prisms and truncated-tetrahedra. Define $\mathbb{P}(\mathcal{C}_{X})=\{\text{vertical edges of }\mathcal{C}_{X}\}\cup\{\text{trapezoids}\}\cup\{\text{triangular blocks}\}\cup\{\text{quadrilateral blocks}\}$. If each component, $\mathbb{P}_{i}$ of $\mathbb{P}(\mathcal{C}_{X})$ is a product $I$ bundle we call $\mathbb{P}(\mathcal{C}_{X})$ a combinatorial product and write $\mathbb{P}_{i}=K_{i}\times I$ where $K_{i}\times 0$ and $K_{i}\times 1$ are isomorphic subcomplexes of $S$. If each $K_{i}$ is simply connected we say $\mathbb{P}(\mathcal{C}_{X})$ is a trivial combinatorial product.
A hexagonal face of a truncated-prism are identified to a hexagonal face of a truncated-tetrahedron or of a (possibly the same) truncated-prism such that the horizontal edges of the faces are identified iff the vertices that their lifts separate in the corresponding tetrahedron $\widetilde{\Delta}$ are identified in $\mathcal{T}$. Define for each hexagonal face of $\mathcal{C}_{X}$ a midpoint and an arc connecting the two midpoints in each truncated-prism. The graph, $\Gamma$, defined by these arcs and midpoints has components homeomorphic to intervals and copies of $S^{1}$. The intervals have endpoints in hexagonal faces of truncated-tetrahedra. If each component $\Gamma$ is homeomorphic to an interval we say $\mathcal{C}_{X}$ is simple. We call the collection of truncated-prisms containing each component of $\Gamma$ a chain.
If $C_{X}$ is simple and $\mathbb{P}(\mathcal{C}_{X})$ is a trivial combinatorial product, we can construct an ideal triangulation of $X$ by crushing the cells of $\mathcal{C}_{X}$ as follows: each component of $S$ is crushed to a point, components of $\mathbb{P}(\mathcal{C}_{X})$ are crushed to edges in the ideal triangulation, truncated-prisms are crushed to faces, and truncated-tetrahedra become ideal tetrahedra.
@@
If one crushes all cells of $\mathcal{C}_{X}$ except for the truncated tetrahedra, the obtained cell-complex is the same cell-decomposition of $M$ described in Section 2.1 used to compute the boundary-slope of spun-normal surfaces. Attaching the cone over the unglued triangular faces of this cell-complex recovers the ideal triangulation $\mathcal{T}^{*}$ above.
@@ 3.2. Inflating Triangulations
Let $\mathcal{T}$ be a triangulation of a 3-manifold $M$. We say $\mathcal{T}$ is a minimal-vertex triangulation if for any triangulation $\mathcal{T}^{\prime}$ of $M$ $|\mathcal{T}^{(0)}|\leq|\mathcal{T}^{\prime(0)}|$. In the case $\partial M$ is homeomorphic to a torus, a minimal-vertex triangulation has exactly one vertex [21]. This induces a 1-vertex triangulation on $\partial M$, that is, the two-triangle triangulation of the torus.
If $\mathcal{T}$ has normal boundary and $\mathcal{T}$ can be crushed along the boundary-linking surface, we say $\mathcal{T}$ can be crushed along $\partial M$. Let $\mathcal{T}^{*}$ be an ideal triangulation of $M$, $\mathcal{T}$ a triangulation of $M$. We say $\mathcal{T}$ is an inflation of $\mathcal{T}^{*}$ if $\mathcal{T}$ has normal boundary and can be crushed along $\partial M$ such that the ideal triangulation obtained by crushing along $M$ is isomorphic to $\mathcal{T}^{*}$. More generally if $\mathcal{T}$ has the above properties we say $\mathcal{T}$ is an inflation triangulation when we do not wish to emphasize $\mathcal{T}^{*}$.
@@
For a surface $S$, a spine, is an embedded 1-complex, such that each component of its complement an open disc. Given a triangulation of $S$, a subcomplex, $\xi$ of the 1-skeleton is called a frame if it is a spine of $S$ and is minimal among all spines of $S$ relative to set inclusion. Notably, the complement of a frame is connected and for the torus there are, up to homeomorphism, two possible frames.
@@
Given an ideal triangulation $\mathcal{T}^{*}$ of $M$, recall in the context of this work $\mathcal{T}^{*}$ has a single ideal vertex. Let $S$ be the vertex-linking surface in $\mathcal{T}^{*}$ and let $\xi$ be a frame of $S$ in the triangulation given by the normal triangles. The inflation begins with the tetrahedra in $\widetilde{\mathbf{\Delta}}^{*}$. Given an edge, $\varepsilon$ of $\xi$ in a 2-simplex, $\sigma$ of $\mathcal{T}^{*}$, we unglue the two tetrahedra at $\sigma$ and attach a new tetrahedron $\widetilde{b}_{\varepsilon}$ called a band tetrahedron as in Figure 2. If there are multiple edges of $\xi$ in $\sigma$ we repeat this process. If for a given edge, $e$ of $\mathcal{T}^{*}$ there are edges of $\xi$ ending at $e$ such that their projection along $e$ onto a regular disc neighborhood of the corresponding vertex in $S$ has a crossing, then we add a new tetrahedron $\widetilde{c}_{e}$ called a crossing tetrahedron. Finally, for each branch point, $v$, we add $index(v)-2$ new tetrahedra $\widetilde{p}_{1},\widetilde{p}_{2}$ called branch-point tetrahedra. Face identifications are determined as in [23] so that $\mathcal{T}_{\xi}=(\widetilde{\mathbf{\Delta}}_{\xi},\Phi)$ is a minimal vertex triangulation of $M$, where $\widetilde{\mathbf{\Delta}}_{\xi}=\widetilde{\mathbf{\Delta}}^{*}\sqcup\{\widetilde{b}_{\varepsilon}\}\sqcup\{\widetilde{c}_{e}\}\sqcup\{\widetilde{p}_{i}\}$. The triangulation $\mathcal{T}_{\xi}$ has normal boundary and admits a crushing along $\partial M$ yielding $\mathcal{T}^{*}$.
@@
The lift of the cells of the boundary-linking surface, $B$, meet the tetrahedra of $\widetilde{\mathbf{\Delta}}_{\xi}$ in a distinguishable way. Let $X\cong M$ be the closure of the component of the complement of $B$ not containing the vertex of $\mathcal{T}_{\xi}$. Each tetrahedron of $\widetilde{\mathbf{\Delta}}_{\xi}$ contains exactly one 3-cell of the induced cell-decomposition $\mathcal{C}_{X}$. Namely, for $\widetilde{\Delta}^{*}\in\widetilde{\mathbf{\Delta}}^{*}$, $\widetilde{\Delta}^{*}$ contains a truncated-tetrahedron; $\widetilde{b}_{\varepsilon}$ contains a truncated-prism; $\widetilde{c}_{e}$ contains a quadrilateral block; $\widetilde{p}_{i}$ contains a triangular block. The chains of truncated-prisms have length at most three. For each edge, $\varepsilon$ in $\xi$ not incident to a branch point, the face of a band tetrahedron $\widetilde{b}_{\varepsilon}$ containing the trapezoidal face of the truncated-prism is glued to another band tetrahedron or some $\widetilde{c}_{e}$ such that the normal arc corresponding to the normal quadrilateral glues to the normal arc of a normal quadrilateral in the to be identified tetrahedron. If $\widetilde{b}_{\varepsilon}$ corresponds to an edge ending at a branch point, the appropriate face containing the trapezoidal face of the truncated-prism glues to a face of some $\widetilde{p}_{i}$ or if necessary some $\widetilde{c}_{e}$ which is in turn glued to $\widetilde{p}_{i}$. Exactly one face of each $\widetilde{p}_{i}$ remains unglued, together forming the material boundary of $\mathcal{T}_{\xi}$.
@@
The above description can be used to identify a given triangulation, $\mathcal{T}=(\widetilde{\mathbf{\Delta}},\mathbf{\Phi})$ of $M$, as the inflation of some ideal triangulation $\mathcal{T}^{*}$ of $M$. In particular, we can write $\widetilde{\mathbf{\Delta}}=\widetilde{\mathbf{\Delta}}^{*}\sqcup\{\widetilde{b}_{i}\}\sqcup\{\widetilde{c}_{j}\}\sqcup\{\widetilde{p}_{k}\}$, where $B,X$ are given as above, $\widetilde{\mathbf{\Delta}}^{*}$ is the collection of tetrahedra containing a truncated-tetrahedron of $\mathcal{C}_{X}$; $\{\widetilde{b}_{i}\}$ the collection containing a truncated-prism; $\{\widetilde{c}_{j}\}$ the collection containing a quadrilateral block; $\{\widetilde{p}_{k}\}$ the collection containing a triangular block. Finally, $\mathcal{C}_{X}$ must be simple; from the inflation procedure a chain can contain at most three truncated prisms.
@@ 3.3. Short Inflations
Let $\mathcal{T}$ be a triangulation of $M$, and $\mathcal{T}^{\prime}$ a triangulation of $M$ obtained after some 2-3 move. If $S$ is a normal surface in $\mathcal{T}$, there exists a normal surface $S^{\prime}$ in $\mathcal{T}^{\prime}$ such that $S$ is isotopic to $S^{\prime}$ as properly embedded surfaces in $M$. Similarly, if the cell-decomposition on $S$ does not include both of the two normal quadrilateral types that separate the diagonal of the base of a square pyramid, then after a 2-2 move on the square pyramid there is again a normal surface $S^{\prime}$ in $\mathcal{T}^{\prime}$ isotopic to $S$. Further if $\Delta,\Delta^{\prime}$ are corresponding tetrahedra in $\mathcal{T},\mathcal{T}^{\prime}$ unaffected by the Pachner move, then the normal discs of $S,S^{\prime}$ in $\Delta,\Delta^{\prime}$ are in 1-1 correspondence. If $S$ meets the bi-pyramid $P$ in a disc formed by a normal quadrilateral as in the left of Figure 3, then after a 2-3 move the disc is divided into two normal quadrilaterals and one normal triangle in $S^{\prime}$. If $S$ meets $P$ in a disc formed by two normal quadrilaterals as in the right of Figure 3, then after a 2-3 move the disc is divided into two normal triangles and one normal quadrilateral in $S^{\prime}$.
@@
By Remark 3.3 any such triangulation $\mathcal{T}^{\prime}$ obtained by a 2-3 move has normal boundary. We wish to consider two particular 2-3 moves: shortening moves, a 2-3 move performed at a face corresponding to a trapezoid in $\mathcal{C}_{X}$ that is common to two band tetrahedra or shared by a crossing tetrahedron and a band tetrahedron; and site-swapping moves (or simply site-swaps), a 2-3 move performed at a face corresponding to a trapezoid common to a band tetrahedron and a branch-point tetrahedron.
@@
(between a band tetrahedron and a crossing tetrahedron) Let $b_{i},c_{j}$ be a band tetrahedron and a crossing tetrahedron respectively, identified at face $\sigma$, the lift of which contains a trapezoidal face of $\mathcal{C}_{X}$. There are two distinct edges, $\tilde{e}_{1}=\langle ab\rangle,\tilde{e}_{2}=\langle cd\rangle$ in the lift $\tilde{c}_{j}$ that lie in the material boundary of $\mathcal{T}$. One of these edges, say $\tilde{e}_{1}$, also lies in $\tilde{b}_{i}$. Consider the lifts of the intersection of $B$ with the bipyramid, $P$ formed from $\tilde{b}_{i},\tilde{c}_{j}$. After performing a 2-3 move at $\sigma$, the unique tetrahedron $\widetilde{\Delta}_{\tilde{e}_{1}}$ containing $\tilde{e}_{1}$, contains a normal quadrilateral of $B^{\prime}$ separating $\tilde{e}_{1}$. For the two vertices of $\widetilde{\Delta}_{\tilde{e}_{1}}$ not in the closure of $\tilde{e}_{1}$ the normal triangle that separates the respective vertex is in $B^{\prime}$. The other two tetrahedra have the edge $\tilde{e}_{2}$ in common. In each, there is a single quadrilateral separating $\tilde{e}_{2}$ and for the appropriate vertices not in the closure of $\tilde{e}_{2}$, a normal triangle separating that vertex. Thus the intersection of each tetrahedra of $B^{\prime}$ with $P$ forms a truncated prism in the induced cell-decomposition $\mathcal{C}^{\prime}_{X}$. As $\mathcal{C}_{X}$ is simple, by Remarks 3.3 and 3.2 the cell-decomposition $\mathcal{C}^{\prime}_{X}$ induced on $X$ by $\mathcal{T}^{\prime}$ is simple. There is a trapezoidal face in the interior of $P$, the closure of which is connected; the preimage of all other components $\mathbb{P}^{\prime}_{i}$ of $\mathbb{P}(\mathcal{C}^{\prime}_{X})$ are a component $\mathbb{P}_{j}$ of $\mathbb{P}(\mathcal{C}_{X})$. Thus $\mathbb{P}(\mathcal{C}^{\prime}_{X})$ is a trivial combinatorial product. By Theorem 3.1, $\mathcal{T}^{\prime}$ is an inflation triangulation of $M$. The number of band tetrahedra $\beta^{\prime}=\beta+2$, the number of crossing tetrahedra $\gamma^{\prime}=\gamma-1$; hence $len(\mathcal{T}^{\prime})=len(\mathcal{T})$.

### v2

@@ 2.2. Fibers and lw-Taut Surfaces
Let $F$ be a compact, orientable, once-punctured surface with $h\in\text{Aut}(F)$. Let $M_{h}$ be the mapping torus $M_{h}=F\times[0,1]/\sim$, where $(x,0)\sim(h(x),1)$. We say a 3-manifold $M$ fibers over the circle $S^{1}$ with fiber $F$ if $M$ is homeomorphic to $M_{h}$ for some $h\in\text{Aut}(F)$. This gives a proper embedding of $F$ in $M$ such that the image is incompressible and $\partial$-incompressible. The identification of $M$ with the mapping torus has strong implications on the homology groups $H_{2}(M)$ and $H_{2}(M,\partial M)$. In particular, any incompressible surface properly embedded in $M$ that is homologous to $F$ is homotopic into $F$ [35]. In the case of knots in rational homology 3-spheres this relation is stronger.
@@
An oriented, incompressible, $\partial$-incompressible surface, $S$, properly embedded in $M$ is said to be taut if the homology class $[S]\in H_{2}(M,\partial M)$ is non-trivial, there is no homologically trivial union of components of $S$, and $S$ is $\chi_{-}$-minimizing over all properly embedded surfaces representing $[S]$. If for a triangulation $\mathcal{T}$, $S$ is transverse to the 2-skeleton of $\mathcal{T}$, we define the weight of $S$ by $wt(S)=|S\cap\mathcal{T}^{(0)}|$. We say $S$ is lw-taut if it has minimal weight among all taut surfaces representing the homology class $[S]$. Define $kS$ to be the disjoint union of $k$ copies of $S$. If $S$ is lw-taut it follows that $kS$ is an lw-taut surface with homology class $k[S]$. The following from [43] gives existence of normal surfaces that are lw-taut:
@@
For a normal surface $S$, we say the carrier $C_{S}$ is taut (lw-taut) if every surface carried by $C_{S}$ is taut (lw-taut). The study of faces of $\mathcal{P}_{\mathcal{T}}$ that carry only minimal weight normal surfaces began with Jaco and Oertel [20] which led to an algorithm to detect if a 3-manifold was Haken (contains an essential surface of genus $\geq$ 1). Again from [43] we have the existence of such faces:
@@ 3.1. Crushing Triangulations along Normal Surfaces
We say a 1- or 2-cell of $\mathcal{C}_{X}$ is horizontal if their interior is in $S$ and vertical otherwise. The quadrilateral vertical 2-cells are called trapezoids. There are two trapezoids in a truncated-prism, three in a triangular block, and four in a rectangular block. All other vertical 2-cells are hexagons found in truncated-prisms and truncated-tetrahedra. Define $\mathbb{P}(\mathcal{C}_{X})=\{\text{vertical edges of }\mathcal{C}_{X}\}\cup\{\text{trapezoids}\}\cup\{\text{triangular blocks}\}\cup\{\text{quadrilateral blocks}\}$. If each component $\mathbb{P}_{i}$ of $\mathbb{P}(\mathcal{C}_{X})$ is a product $I$ bundle, then we call $\mathbb{P}(\mathcal{C}_{X})$ a combinatorial product and write $\mathbb{P}_{i}=K_{i}\times I$ where $K_{i}\times 0$ and $K_{i}\times 1$ are isomorphic subcomplexes of $S$. If each $K_{i}$ is simply connected we say $\mathbb{P}(\mathcal{C}_{X})$ is a trivial combinatorial product.
A hexagonal face of a truncated-prism is identified to a hexagonal face of a truncated-tetrahedron or of a (possibly the same) truncated-prism such that the horizontal edges of the faces are identified if and only if the vertices that their lifts separate in the corresponding tetrahedron $\widetilde{\Delta}$ are identified in $\mathcal{T}$. Define for each hexagonal face of $\mathcal{C}_{X}$ a midpoint and an arc connecting the two midpoints in each truncated-prism. The graph $\Gamma$ defined by these arcs and midpoints has components homeomorphic to intervals and copies of $S^{1}$. The intervals have endpoints in hexagonal faces of truncated-tetrahedra. If each component $\Gamma$ is homeomorphic to an interval we say $\mathcal{C}_{X}$ is simple. We call the collection of truncated-prisms containing each component of $\Gamma$ a chain.
If $C_{X}$ is simple and $\mathbb{P}(\mathcal{C}_{X})$ is a trivial combinatorial product, we can construct an ideal triangulation of $X$ by crushing the cells of $\mathcal{C}_{X}$ as follows: each component of $S$ is crushed to a point, components of $\mathbb{P}(\mathcal{C}_{X})$ are crushed to edges in the ideal triangulation, truncated-prisms are crushed to faces, and truncated-tetrahedra become ideal tetrahedra, see Figure 2.
Figure 2. Shown are the results of crushing each cell type of $\mathcal{C}_{X}$. The gray faces represent cells of the normal surface, the rounded vertices represent the ideal vertex of $\mathcal{T}^{*}$.
@@
If one crushes all cells of $\mathcal{C}_{X}$ except for the truncated tetrahedra, the resulting cell-complex is the same cell-decomposition of $X$ described in Section 2.1 used to compute the boundary-slope of spun-normal surfaces. Attaching the cone over the unglued triangular faces of this cell-complex recovers the ideal triangulation $\mathcal{T}^{*}$ of $X$ in Theorem 3.1 above.
@@ 3.2. Inflating Triangulations
Let $\mathcal{T}$ be a triangulation of a 3-manifold $M$. We say $\mathcal{T}$ is a minimal-vertex triangulation if for any triangulation $\mathcal{T}^{\prime}$ of $M$ $|\mathcal{T}^{(0)}|\leq|\mathcal{T}^{\prime(0)}|$. If $\partial M$ is connected, then a minimal-vertex triangulation has exactly one vertex [21], inducing a 1-vertex triangulation on $\partial M$; when $\partial M$ is homeomorphic to a torus, this is the standard two-triangle triangulation of the torus.
If $\mathcal{T}$ has normal boundary and can be crushed along the boundary-linking surface, we say $\mathcal{T}$ can be crushed along $\partial M$. Let $\mathcal{T}^{*}$ be an ideal triangulation of $M$, $\mathcal{T}$ a triangulation of $M$. We say $\mathcal{T}$ is an inflation of $\mathcal{T}^{*}$ if $\mathcal{T}$ has normal boundary and can be crushed along $\partial M$ such that the ideal triangulation obtained by crushing along $M$ is isomorphic to $\mathcal{T}^{*}$. More generally if $\mathcal{T}$ has the above properties we say $\mathcal{T}$ is an inflation triangulation when we do not wish to emphasize $\mathcal{T}^{*}$.
@@
For a surface $S$, a spine is an embedded 1-complex such that each component of its complement is an open disc. Given a triangulation of $S$, a subcomplex $\xi$ of the 1-skeleton is called a frame if it is a spine of $S$ and is minimal among all spines of $S$ relative to set inclusion. Notably, the complement of a frame is connected and for the torus there are, up to homeomorphism, two possible frames.
@@
Given an ideal triangulation $\mathcal{T}^{*}$ of $M$, recall in the context of this work $\mathcal{T}^{*}$ has a single ideal vertex. Let $S$ be the vertex-linking surface in $\mathcal{T}^{*}$ and let $\xi$ be a frame of $S$ in the triangulation given by the normal triangles. The inflation begins with the tetrahedra in $\widetilde{\mathbf{\Delta}}^{*}$. Given an edge $\varepsilon$ of $\xi$ in a 2-simplex $\sigma$ of $\mathcal{T}^{*}$, we unglue the two tetrahedra at $\sigma$ and attach a new tetrahedron $\widetilde{b}_{\varepsilon}$ called a band tetrahedron as in Figure 3. If there are multiple edges of $\xi$ in $\sigma$ we repeat this process. If for a given edge, $e$ of $\mathcal{T}^{*}$ there are edges of $\xi$ ending at $e$ such that their projection along $e$ onto a regular disc neighborhood of the corresponding vertex in $S$ has a crossing, then we add a new tetrahedron $\widetilde{c}_{e}$ called a crossing tetrahedron. Finally, for each branch point, $v$, we add $index(v)-2$ new tetrahedra $\widetilde{p}_{1},\widetilde{p}_{2}$ called branch-point tetrahedra. Face identifications are determined as in [23] so that $\mathcal{T}_{\xi}=(\widetilde{\mathbf{\Delta}}_{\xi},\Phi)$ is a minimal vertex triangulation of $M$, where $\widetilde{\mathbf{\Delta}}_{\xi}=\widetilde{\mathbf{\Delta}}^{*}\sqcup\{\widetilde{b}_{\varepsilon}\}\sqcup\{\widetilde{c}_{e}\}\sqcup\{\widetilde{p}_{i}\}$. The triangulation $\mathcal{T}_{\xi}$ has normal boundary and admits a crushing along $\partial M$ yielding $\mathcal{T}^{*}$.
@@
The lift of the cells of the boundary-linking surface $B$ meet the tetrahedra of $\widetilde{\mathbf{\Delta}}_{\xi}$ in a distinguishable way. Let $X\cong M$ be the closure of the component of the complement of $B$ not containing the vertex of $\mathcal{T}_{\xi}$. Each tetrahedron of $\widetilde{\mathbf{\Delta}}_{\xi}$ contains exactly one 3-cell of the induced cell-decomposition $\mathcal{C}_{X}$. Namely, for $\widetilde{\Delta}^{*}\in\widetilde{\mathbf{\Delta}}^{*}$, $\widetilde{\Delta}^{*}$ contains a truncated-tetrahedron, $\widetilde{b}_{\varepsilon}$ contains a truncated-prism, $\widetilde{c}_{e}$ contains a quadrilateral block, and $\widetilde{p}_{i}$ contains a triangular block. The chains of truncated-prisms have length at most three. For each edge $\varepsilon$ in $\xi$ not incident to a branch point, the face of a band tetrahedron $\widetilde{b}_{\varepsilon}$ containing the trapezoidal face of the truncated-prism is glued to another band tetrahedron or some $\widetilde{c}_{e}$ such that the normal arc corresponding to the normal quadrilateral glues to the normal arc of a normal quadrilateral in the to be identified tetrahedron. If $\widetilde{b}_{\varepsilon}$ corresponds to an edge ending at a branch point, the appropriate face containing the trapezoidal face of the truncated-prism glues to a face of some $\widetilde{p}_{i}$ or if necessary some $\widetilde{c}_{e}$ which is in turn glued to $\widetilde{p}_{i}$. Exactly one face of each $\widetilde{p}_{i}$ remains unglued, together forming the material boundary of $\mathcal{T}_{\xi}$.
@@
The above description can be used to identify a given triangulation $\mathcal{T}=(\widetilde{\mathbf{\Delta}},\mathbf{\Phi})$ of $M$ as the inflation of some ideal triangulation $\mathcal{T}^{*}$ of $M$. In particular, we can write $\widetilde{\mathbf{\Delta}}=\widetilde{\mathbf{\Delta}}^{*}\sqcup\{\widetilde{b}_{i}\}\sqcup\{\widetilde{c}_{j}\}\sqcup\{\widetilde{p}_{k}\}$, where $B,X$ are given as above. Here $\widetilde{\mathbf{\Delta}}^{*}$ is the collection of tetrahedra containing a truncated-tetrahedron of $\mathcal{C}_{X}$, $\{\widetilde{b}_{i}\}$ the collection containing a truncated-prism, $\{\widetilde{c}_{j}\}$ the collection containing a quadrilateral block, $\{\widetilde{p}_{k}\}$ the collection containing a triangular block. Finally, $\mathcal{C}_{X}$ must be simple; from the inflation procedure a chain can contain at most three truncated prisms.
@@ 3.3. Short Inflations
Let $\mathcal{T}$ be a triangulation of $M$, and $\mathcal{T}^{\prime}$ a triangulation of $M$ obtained after some 2-3 move on $\mathcal{T}$. If $S$ is a normal surface in $\mathcal{T}$, then there exists a normal surface $S^{\prime}$ in $\mathcal{T}^{\prime}$ such that $S$ is isotopic to $S^{\prime}$ as properly embedded surfaces in $M$. Similarly, if the cell-decomposition on $S$ does not include both of the two normal quadrilateral types that separate the diagonal of the base of a square pyramid, then after a 2-2 move on the square pyramid there is again a normal surface $S^{\prime}$ in $\mathcal{T}^{\prime}$ isotopic to $S$. Further if $\Delta,\Delta^{\prime}$ are corresponding tetrahedra in $\mathcal{T},\mathcal{T}^{\prime}$ unaffected by the Pachner move, then the normal discs of $S,S^{\prime}$ in $\Delta,\Delta^{\prime}$ are in 1-1 correspondence. If $S$ meets the bi-pyramid $P$ in a disc formed by a normal quadrilateral as in the left of Figure 4, then after a 2-3 move the disc is divided into two normal quadrilaterals and one normal triangle in $S^{\prime}$. If $S$ meets $P$ in a disc formed by two normal quadrilaterals as in the right of Figure 4, then after a 2-3 move the disc is divided into two normal triangles and one normal quadrilateral in $S^{\prime}$.
@@
By Remark 3.3 any such triangulation $\mathcal{T}^{\prime}$ obtained by a 2-3 move on an inflation triangulation has normal boundary. We wish to consider two particular 2-3 moves: shortening moves, a 2-3 move performed at a face corresponding to a trapezoid in $\mathcal{C}_{X}$ that is common to two band tetrahedra or shared by a crossing tetrahedron and a band tetrahedron; and site-swapping moves (or simply site-swaps), a 2-3 move performed at a face corresponding to a trapezoid common to a band tetrahedron and a branch-point tetrahedron.
@@
(between a band tetrahedron and a crossing tetrahedron) Let $b_{i},c_{j}$ be a band tetrahedron and a crossing tetrahedron respectively, identified at face $\sigma$, the lift of which contains a trapezoidal face of $\mathcal{C}_{X}$. There are two distinct edges $\tilde{e}_{1}=\langle ab\rangle,\tilde{e}_{2}=\langle cd\rangle$ in the lift $\tilde{c}_{j}$ that lie in the material boundary of $\mathcal{T}$. One of these edges, say $\tilde{e}_{1}$, also lies in $\tilde{b}_{i}$. Consider the lifts of the intersection of $B$ with the bipyramid $P$ formed from $\tilde{b}_{i},\tilde{c}_{j}$. After performing a 2-3 move at $\sigma$, the unique tetrahedron $\widetilde{\Delta}_{\tilde{e}_{1}}$ containing $\tilde{e}_{1}$ contains a normal quadrilateral of $B^{\prime}$ separating $\tilde{e}_{1}$. For the two vertices of $\widetilde{\Delta}_{\tilde{e}_{1}}$ not in the closure of $\tilde{e}_{1}$ the normal triangle that separates the respective vertex is in $B^{\prime}$. The other two tetrahedra have the edge $\tilde{e}_{2}$ in common. In each, there is a single quadrilateral separating $\tilde{e}_{2}$, and for the appropriate vertices not in the closure of $\tilde{e}_{2}$ a normal triangle separating that vertex. Thus the intersection of each tetrahedra of $B^{\prime}$ with $P$ forms a truncated prism in the induced cell-decomposition $\mathcal{C}^{\prime}_{X}$. As $\mathcal{C}_{X}$ is simple, by Remarks 3.3 and 3.2 the cell-decomposition $\mathcal{C}^{\prime}_{X}$ induced on $X$ by $\mathcal{T}^{\prime}$ is simple. There is a trapezoidal face in the interior of $P$, the closure of which is connected; the preimage of all other components $\mathbb{P}^{\prime}_{i}$ of $\mathbb{P}(\mathcal{C}^{\prime}_{X})$ are a component $\mathbb{P}_{j}$ of $\mathbb{P}(\mathcal{C}_{X})$. Thus $\mathbb{P}(\mathcal{C}^{\prime}_{X})$ is a trivial combinatorial product. By Theorem 3.1, $\mathcal{T}^{\prime}$ is an inflation triangulation of $M$. The number of band tetrahedra $\beta^{\prime}=\beta+2$, the number of crossing tetrahedra $\gamma^{\prime}=\gamma-1$; hence $len(\mathcal{T}^{\prime})=len(\mathcal{T})$.